### v1

# A Regret Perspective on Online Selective Generation

Abstract
Large language generative models increasingly interact with humans, while their falsified responses raise concerns. To address this hallucination effect, selectively abstaining from answering, called selective generation, provides an effective way for generators to control the hallucination when it is unsure of their answers. However, as selective generators are interacting under non-stochastic environments and having partial feedback from users on selective generation (e.g., thumbs up or down on the selected answer), learning methods for selective generation under such practical setups are crucial but currently missing. To address these limitations, we propose an online learning algorithm for selective generation under partial feedback. In particular, as learning under partial feedback is well-studied by multi-armed bandit problems, we reduce selective generation to bandits and provide a novel conversion lemma from bandits back to selective generation to leverage any known bandit algorithms and theoretical properties. This mainly connects regret guarantees of bandits to false discovery rate (FDR) guarantees of selective generation for controlling hallucination. However, naively exploiting known bandit algorithms and their regret bounds suffers from slow convergence speed in practice due the nature of partial feedback. To overcome this, we exploit a unique structure of arms in selective generation for feedback unlocking, i.e., unlocking unknown feedback from observed feedback. We theoretically and empirically evaluate the efficacy of the proposed online selective generation algorithm under partial feedback over diverse data environment setups, resulting in controlling a desired FDR, while maintaining reasonable selection efficiency, i.e., the ratio of non-abstaining answers, compared to baselines.
{NoHyper}**footnotetext: Equal contribution
1 Introduction
As large language generators openai2023gpt35turbo ; meta2024llama31-8binstruct surpass average human performance, their interaction with humans becomes increasingly prevalent. However, as they tend to generate incorrect information, or so-called hallucinated information, community’s concerns on their mis-alignment issues have significantly increased. One simple but effective approach to mitigating hallucination is selective generation, selectively abstaining from generation if the generator is unsure of its response geifman2017selective ; goren2024hierarchical ; mohri2023learning ; lee2024selective . By the abstention, the selective generator may control its high precision, or equivalently a low false discovery rate (FDR), i.e., whenever it says answers, they are mostly correct. Interestingly, few selective prediction and generation methods geifman2017selective ; goren2024hierarchical ; lee2024selective provide theoretical guarantees on the controllability of the FDR to ensure the trustworthiness of generators.
However, conventional selective prediction methods geifman2017selective ; goren2024hierarchical ; lee2024selective are designed under limited stochastic assumptions, i.e., data are independently drawn from a fixed distribution, which undermines their applicability to real world applications in adversarial and distribution-shifted environments. Moreover, the methods require full feedback (e.g., a true answer is given) instead of partial feedback (e.g., the correctness of a generated answer), where partial feedback is more practical and easier to obtain (e.g., thumbs-up buttons in dialog systems).
To mitigate these limitations, we propose a novel online selective generation algorithm to control a false discovery rate (FDR), while maximizing selection efficiency (i.e., the ratio of non-abstaining cases), with partial feedback. First, we address the online learning problem under partial feedback by leveraging profound achievements from online learning and multi-armed bandits. To this end, we reduce selective generation to bandits, exploit any regret minimization algorithms with regret bounds, and then translate the regret bounds back into FDR bounds for selective generation by introducing a novel Regret-to-FDR conversion lemma. This enables us to use any regret minimization algorithms for selective generation with FDR controllability guarantees.
Second, we design a regret minimization algorithm tailored for selective generation to fully exploit feedback information. In particular, partial feedback is practical but the amount of new information is clearly insufficient compared to full feedback, suffering from sample inefficiency and slower convergence speed. To address this technical challenge, we exploit a unique structure of selective generators to unlock additional feedback from observed partial feedback, dubbed by partial feedback unlocking. By exploiting this, we extend the Exp3 algorithm auer2002nonstochastic for adversarial bandits for online Selective generation with partial feedback UnLocking (ExSUL). Moreover, we provide the $\mathcal{{O}}(\ell_{\text{max}}\sqrt{T\ln|\mathcal{{H}}|})$ regret bound of ExSUL, which is better than a regret minimizer with partial feedback, i.e., $\mathcal{{O}}(\ell_{\text{max}}\sqrt{T|\mathcal{{H}}|\ln|\mathcal{{H}}|})$, and compatible to a regret minimizer with full feedback, i.e., $\mathcal{{O}}(\ell_{\text{max}}\sqrt{T\ln|\mathcal{{H}}|})$, where $T$ is a time horizon and $\mathcal{{H}}$ is a set of hypotheses of selective generators. See Figure 1 for qualitative results of our method that controls an FDR in an interactive environment.
We empirically evaluate the efficacy of the proposed online selective generation algorithm ExSUL. In particular, we consider two tasks (including question-answering over TriviaQA joshi2017triviaqa and Natural Question kwiatkowski2019natural and dialog conversations via two dialog agents), three learning environments (including stochastic, distribution-shifted, and interactive environments), and two language models (including GPT-3.5-turbo openai2023gpt35turbo and LLaMA3.1 meta2024llama31-8binstruct ) to show that our method (1) controls a desired FDR and (2) maintains reasonably low selection inefficiency, i.e., the ratio of abstention. Additionally, our learner’s FDR convergence speed is clearly better than a regret minimizer with partial feedback and compatible to a regret minimizer with full feedback (which forms a performance upper bound).
Figure 1: Qualitative examples from interactive dialog simulation. This demonstrates that our proposed method ExSUL effectively controls the rate of hallucination in the FDR by abstaining from answering under a practical online setup with partial feedback. See Section 5.3 for details.
1.1 Related Work
Here, we introduce tightly related literature, including selective prediction, online learning, and bandit problems. See Appendix A for additional related work.
Selective Prediction. Selective prediction is a method that abstains from making a prediction to control the FDR in a certified manner. geifman2017selective proposes selective classification that uses a threshold for a scoring function to abstain from uncertain predictions. goren2024hierarchical extends this to exploit hierarchical structure of labels for better efficiency. mohri2023learning extends this idea to generation tasks but mainly focuses on decontextualization task without the FDR guarantee. lee2024selective further generalizes selective prediction to generation tasks by introducing the concept of an entailment set, which addresses the challenge of open-ended question-answering. In particular, they propose a semi-supervised method to leverage unlabeled entailment data, providing a theoretical guarantee for controlling the FDR at a desired level. Previous work focuses on a batch learning setup under a stochastic assumption, which may be fragile in real-world applications. Contrast to this, we consider a non-stochastic setup, allowing distribution shift, to propose online selective generation methods.
Online Learning and Bandit Problems. Sequential prediction cesa2006prediction ; mohri2012foundations ; sankararaman2016semi ; foster2023foundations designs learners that adapt to sequentially arriving data. Online learning, e.g., exponential weighting littlestone1994weighted , mostly assumes adversarial environments where data can be arbitrarily, possibly adaptively, manipulated by adversaries. Under this environment, a learner receives full feedback on predictions. In contrast to this, bandit problems, including multi-armed bandits agrawal1995sample and adversarial bandits auer2002nonstochastic , assume to have sequentially-arriving partial feedback, i.e., a learner receives feedback on a chosen arm over time. To efficiently leverage partial feedback, structured bandits russo2013eluder exploit the functional structure of arms-to-loss functions and semi-bandits sankararaman2016semi allow to choose a set of arms with fixed size for better learning efficiency. We exploit online adversarial bandits to mitigate stochastic assumptions in learning traditional selective prediction and to consider practical learning under partial feedback. Moreover, we leverage a specific structure between arms and loss in selective generation for better learning efficiency.
2 Preliminary
We introduce preliminaries on language generation, selective prediction, and adversarial bandits. See Appendix B for online learning. To this end, let $\mathcal{{X}}$ be a set of inputs (e.g., examples or questions) and $\mathcal{{Y}}$ be a set of outputs (e.g., labels or answers).
2.1 Language Generation
We mainly consider language generators as our model to control hallucination. In particular, let $G:\mathcal{{X}}\to\mathcal{{Y}}$ be a language generator, where $\mathcal{{W}}$ is a set of tokens and $\mathcal{{X}}=\mathcal{{Y}}\coloneqq\cup_{i=0}^{\infty}\mathcal{{W}}^{i}$. Here, each $i$-th token $\hat{\mathbf{y}}_{i}$ of a generated answer $\hat{\mathbf{y}}\in\mathcal{{Y}}$ is decoded from an underlying probability distribution $p(\mathbf{y}\mid\mathbf{x})$, where $p$ is usually learned via language data. For the decoding strategy, we consider greedy decoding, i.e., $\hat{\mathbf{y}}_{i}=\arg\max_{w\in\mathcal{{W}}}p(w\mid\mathbf{x},\hat{\mathbf{y}}_{1:i-1})$, where $\mathbf{y}_{a:b}\coloneqq(\mathbf{y}_{a},\dots,\mathbf{y}_{b})$. Given a generated answer $\hat{\mathbf{y}}$, there are multiple ways to measure its likelihood of correctness $f:\mathcal{{X}}\times\mathcal{{Y}}\to\mathbb{R}$. A standard way $f_{\texttt{std}}$ considers a length-normalized token probability, i.e., $f_{\texttt{std}}(\mathbf{x},\hat{\mathbf{y}})\coloneqq(p(\hat{\mathbf{y}}_{1}\mid\mathbf{x})\prod_{i=2}^{|\hat{\mathbf{y}}|}p(\hat{\mathbf{y}}_{i}\mid\mathbf{x},\mathbf{\hat{y}}_{1:i-1}))^{\nicefrac{{1}}{{|\hat{\mathbf{y}}|}}}$. A better alternative is to consider consistency of $\hat{\mathbf{y}}$ to multiple generated answers with sampling manakul2023selfcheckgpt , i.e., $f_{\texttt{con}}(\mathbf{x},\hat{\mathbf{y}})=\hat{\mathbbm{E}}_{\mathbf{y}^{\prime}\sim G(x)}f_{E}(\mathbf{y}^{\prime},\hat{\mathbf{y}})$, denoted by a consistency score, where we use sampling for decoding $\mathbf{y}^{\prime}$, $f_{E}(\mathbf{y}^{\prime},\hat{\mathbf{y}})$ measures an entailment score of $\hat{\mathbf{y}}$ given $\mathbf{y}^{\prime}$ (i.e., whether $\mathbf{y}^{\prime}$ entails $\hat{\mathbf{y}}$) via an entailment model (e.g., GPT-3.5-turbo).
2.2 Selective Prediction
Selective prediction geifman2017selective and generation lee2024selective provide certified control over the risk of incorrect predictions by abstaining answers (saying IDK) when the prediction is uncertain. Given a predictor $\hat{\mathbf{y}}:\mathcal{{X}}\to\mathcal{{Y}}$, a selective predictor $\hat{{S}}:\mathcal{{X}}\to\mathcal{{Y}}\cup\{\texttt{IDK}\}$ abstains from returning $\hat{\mathbf{y}}(\mathbf{x})$ if a selection function $\hat{{s}}:\mathcal{{X}}\times\mathcal{{Y}}\to\{0,1\}$ deems the prediction uncertain, i.e., $\hat{{S}}(\mathbf{x})\coloneqq\begin{cases}\hat{\mathbf{y}}(\mathbf{x})&\text{if~{}}\hat{{s}}(\mathbf{x},\hat{\mathbf{y}}(\mathbf{x}))=1\\ \texttt{IDK}&\text{otherwise}\end{cases}.$ In learning, the selection function $\hat{{s}}$ is chosen to possibly satisfy a desired level of a false discovery rate, i.e., $\mathbbm{P}(\hat{{S}}(\mathbf{x})\neq_{E}\mathbf{y}\mid\hat{{S}}(\mathbf{x})\neq\texttt{IDK})$ over the i.i.d. samples of $(\mathbf{x},\mathbf{y})$. The equality relation $\neq_{E}$ is usually the standard exact matching in selective classification geifman2017selective , which is extended to capture semantic correctness via entailment in selective generation lee2024selective . In contrast to the standard stochastic setup as in the previous literature, we consider online learning under partial feedback.
2.3 Regret Minimization
In sequential prediction, the main goal is to find a learner whose performance is as good as the best learner in hindsight. As a goodness metric, we use widely accepted expected regret, simply called regret in this paper, leading to regret minimization. The goal of regret minimization is to find a distribution $p_{t}$ over hypotheses $\mathcal{{H}}$ to minimize regret, i.e., the gap between the learner’s expected cumulative loss over a possibly randomly chosen hypothesis $\tau_{t}\sim p_{t}$ and the best cumulative loss for any loss sequence $\ell_{t}$ over a time horizon $T$:
$\textbf{Reg}_{T}\coloneqq\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t})-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau).$ (1)
Here, the best cumulative loss assumes to have access to loss for any $\tau$, i.e., full feedback, but the learner’s cumulative loss may have limited access to loss, e.g., partial feedback. Depending on this feedback setup, we consider two learning problems: online learning with full feedback and adversarial bandits with partial feedback. In the following section, we review adversarial bandits with one regret minimization algorithm, called Exp3. A well-known regret minimization algorithm for online learning with feedback is Exponential Weighting (EW) littlestone1994weighted , introduced in Appendix B.
2.4 Adversarial Bandits and Exp3 Algorithm
Traditional online learning typically assumes full feedback, i.e., loss for any hypothesis is computable. However, practical applications may only have partial feedback, i.e., loss for a chosen hypothesis is only given. To consider learning in partial feedback, we consider multi-armed bandit in an adversarial setting, called adversarial bandits. In adversarial bandits, we consider the same regret as in online learning, except that a learner only receives feedback of a chosen arm by itself, i.e., partial feedback, and an arbitrary loss sequence $\ell_{t}(\tau_{t})$ over a time horizon $T$ is fixed before learning by an oblivious adversary. In bandit context, we denote a hypothesis as an arm.
To minimize the regret $\textbf{Reg}_{T}$, Exponential-weight algorithm for Exploration and Exploitation (Exp3) auer2002nonstochastic is one standard algorithm, which leverages the EW algorithm for adversarial bandits (Algorithm 3). In particular, Exp3 exploits an unbiased estimator ~$\tilde{\ell}_{t}(\tau\mid\tau_{t})$ of loss for all arms $\tau\in\mathcal{{H}}$ constructed from loss $\ell_{t}(\tau_{t})$ of chosen arms $\tau_{t}$, i.e.,
~$\tilde{\ell}_{t}(\tau\mid\tau_{t})\coloneqq{\ell_{t}(\tau)}/{p_{t}(\tau)}\cdot\mathbbm{1}(\tau_{t}=\tau)\in[0,\infty).$ (2)
Note that ~$\tilde{\ell}_{t}(\tau\mid\tau_{t})$ is an unbiased estimator, i.e., $\mathbbm{E}_{\tau_{t}\sim p_{t}}\tilde{\ell}_{t}(\tau\mid\tau_{t})=\ell_{t}(\tau)$; thus, by using this in $\textbf{Reg}_{T}$, we have the following regret bound for Exp3. See Appendix G.2 for a proof.
Theorem 1.
auer2002nonstochastic ; foster2023foundations Let $\ell_{t}(\cdot)\in[0,\ell_{\text{max}}]$. For any $T\in\mathbb{N}$ and finite hypotheses $\mathcal{{H}}$, Algorithm 3 provides the following regret bound if $\eta=\sqrt{{2\ln|\mathcal{H}|}/{(\ell_{\max}^{2}T|\mathcal{H}|)}}$:
$\!\textbf{Reg}_{T}\!=\!\sum_{t=1}^{T}\mathbbm{E}_{\tau,\tau_{t}\sim p_{t}}\tilde{\ell}_{t}(\tau\mid\tau_{t})\!-\!\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau)\!\leq\!\ell_{\max}\sqrt{2T|\mathcal{H}|\ln|\mathcal{H}|}\!=\!\mathcal{{O}}(\ell_{\text{max}}\sqrt{T|\mathcal{{H}}|\ln|\mathcal{{H}}|}).$ (3)
We will exploit Exp3 to extend this for learning selective generators with partial feedback.
3 Problem: Online Selective Generation with Partial Feedback
We consider online learning of a selective generator under partial feedback for language models. In particular, let $\mathcal{{W}}$ be a set of tokens and $\mathcal{{X}}=\mathcal{{Y}}\coloneqq\cup_{i=0}^{\infty}\mathcal{{W}}$ be a set of (token) sequences. Here, given a generator $G:\mathcal{{X}}\to\mathcal{{Y}}$, we consider a time-varying selective generator $\hat{{S}}_{t}:\mathcal{{X}}\to\mathcal{{Y}}\cup\{\texttt{IDK}\}$ that abstains from answering if a generated answer $G(\mathbf{x}_{t})$ at time $t$ is uncertain, i.e., $\hat{{S}}_{t}(\mathbf{x}_{t})\coloneqq\begin{cases}G(\mathbf{x}_{t})&\text{if~{}}\hat{{s}}(\mathbf{x}_{t},G(\mathbf{x}_{t}))=1\\ \texttt{IDK}&\text{otherwise}\end{cases},$ where $\hat{{s}}:\mathcal{{X}}\times\mathcal{{Y}}\to\{0,1\}$ is a selection function and IDK represents “I don’t know”. To learn the selective generator, we consider online learning under a non-stochastic assumption: each time step $t$ until a time horizon $T$, a learning algorithm (1) observes $\mathbf{x}_{t}\in\mathcal{{X}}$, (2) predicts $\hat{{S}}_{t}(\mathbf{x}_{t})$ where $\hat{{S}}_{t}$ is drawn from a learned distribution $p_{t}$ over selective generators, (3) observes partial feedback $b_{t}\coloneqq\mathbbm{1}(\hat{{S}}_{t}(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t})$ where $\mathbf{x}_{t}$ and $\mathbf{y}_{t}$ are drawn from a time-varying distribution, and (4) update the distribution $p_{t}$ over selective generators by using partial feedback. Here, $A\neq_{E}B$ means $A$ and $B$ are different in terms of a given correctness relation $E$, e.g., textual-entailment bowman2015large and $b_{t}=1$ if $\hat{{S}}_{t}(\mathbf{x}_{t})=\texttt{IDK}$ by the definition. Importantly, we consider partial feedback, i.e., we cannot directly observe $\mathbf{y}_{t}$ but we instead observe whether our prediction is correct or not, i.e., $b_{t}$. This is an ordinary setup in using language generators where a user provides thumb up or down feedback if a generated answer is correct or not, respectively.
Under learning with partial feedback, we aim to update a distribution $p_{t}$ over selective generators that controls an empirical false discovery rate (FDR), with respect to a correctness relation $E$ at a desired level $\alpha\in[0,1]$ up to a time horizon $T$, i.e.,
$\textbf{FDR}_{T}\coloneqq\frac{\sum_{t=1}^{T}\mathbbm{E}_{\hat{{S}}_{t}\sim p_{t}}\mathbbm{1}\left(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK}~{}\wedge~{}\hat{{S}}_{t}(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t}\right)}{\sum_{t=1}^{T}\mathbbm{E}_{\hat{{S}}_{t}\sim p_{t}}\mathbbm{1}\left(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK}\right)}\leq\alpha,$ (4)
where $\textbf{FDR}_{T}=\alpha$ if $\hat{{S}}_{t}(\mathbf{x}_{t})=\texttt{IDK}$ for all $t$ to penalize inefficient cases. Equivalently, we wish to find $\hat{{S}}_{1:T}$ that make the following FDR risk smaller than zero: $\sum_{t=1}^{T}\big{[}\mathbbm{1}(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK}\wedge\hat{{S}}_{t}(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t})-\alpha\mathbbm{1}(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK})\big{]}.$ Then, the goal is to learn a distribution $p_{t}$ over selective generators such that drawn selective generator $\hat{{S}}_{t}\sim p_{t}$ controls the FDR at a desired level $\alpha$ for any sequences, i.e.,
$\frac{1}{T}\sum_{t=1}^{T}\mathbbm{E}_{\hat{{S}}_{t}\sim p_{t}}\left[\mathbbm{1}\left(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK}\wedge\hat{{S}}_{t}(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t}\right)-\alpha\,\mathbbm{1}\left(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK}\right)\right]\leq\varepsilon(T)$ (5)
for any $\mathbf{x}_{t}$ and $\mathbf{y}_{t}$, where $\varepsilon(T)$ is some decreasing function in $T$. Additionally, we wish to minimize selection inefficiency $\textbf{Ineff}_{T}\coloneqq\frac{1}{T}\sum_{t=1}^{T}\mathbbm{E}_{\hat{{S}}_{t}\sim p_{t}}\mathbbm{1}(\hat{{S}}_{t}(\mathbf{x}_{t})=\texttt{IDK}).$
4 Method: Online Selective Generation with Partial Feedback Unlocking
We leverage a regret perspective from online learning and bandit problems in designing a learning algorithm for selective generation with full or partial feedback to control the FDR. In particular, we reduce selective generation to online learning with full feedback or bandit problems with partial feedback to use existing regret minimization algorithms. The main benefit of this reduction is that this enables us to leverage any regret minimization algorithms (e.g., EW or Exp3) and their regret bounds for selective generation algorithms with FDR bounds. To connect regret bounds and FDR bounds, we introduce a novel conversion lemma from the regret to the FDR in Section 4.4. However, under partial feedback setups, simply leveraging existing bandit algorithms may not lead to sample efficient algorithms without exploiting properties of selective generation.
In this section, we propose a novel online selective generation with partial feedback unlocking for sample efficiency. First, we introduce the reduction of the online selective generation to adversarial bandits in Section 4.1. In the following Section 4.2, we extend the Exp3 algorithm auer2002nonstochastic for adversarial bandits to learn selective generators by exploiting unique structure in selection functions by feedback unlocking, followed by its novel regret bound in Section 4.3. This regret bound leads to an FDR bound by using our conversion lemma in Section 4.4. Note that online selective generation under full feedback can be devised in a similar way. See Appendix C for this detail via EW.
Notations. We first introduce our choice of the selection function and simplified notations. As in conventional selective prediction literature, we consider the scalar parameterization of a selection function given a scoring function $f:\mathcal{{X}}\times\mathcal{{Y}}\to{[0,1)}$, i.e., $\hat{{s}}(\mathbf{x}_{t},G(\mathbf{x}_{t}))\coloneqq\mathbbm{1}\left(f(\mathbf{x}_{t},G(\mathbf{x}_{t}))\geq\tau\right),$ thus the selective generator $\hat{{S}}$ is parameterized by $\tau$, denoted by $\hat{{S}}(\cdot;\tau)$. Here, the scoring function $f$ measures the likelihood of $G(\mathbf{x}_{t})$ for being an answer of $\mathbf{x}_{t}$. Also, we specifically consider that $\tau$ is chosen from the finely-quantized, finite space of $[0,1]$, denoted by $\mathcal{{H}}$ (i.e., $\tau\in\mathcal{{H}}$). Given this parameterization, we consider the following shorthand to simplify key notations: $a_{t}(\tau)\coloneqq\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau)=\texttt{IDK})$, which measures the selection inefficiency at time $t$, called inefficiency loss, and $d_{t}(\tau,\alpha)\coloneqq\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau)\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau)\neq_{E}\mathbf{y}_{t})-\alpha\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau)\neq\texttt{IDK})+\alpha$, which measures the violation of the FDR constraint at time $t$ with a margin by $\alpha$ to penalize the IDK response, called FDR loss with a margin. Based on these short hands, we define loss for online learning and bandits that simultaneously considers the FDR constraint and selection inefficiency as follows:
$\ell_{t}(\tau,\alpha)\coloneqq a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)\in\{0,\lambda,1+\lambda\alpha\},$ (6)
where $\lambda$ is a regularized parameter, which will be determined via theoretical analysis.
4.1 Reduction: From Online Selective Generation to Adversarial Bandits
Table 1: From online selective generation to adversarial bandits
online selective generation adversarial bandits
models selective generators $\mathcal{{H}}$ finite arms $\mathcal{{H}}$
feedback $\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq_{E}\mathbf{y}_{t})$ $\ell_{t}(\tau_{t},\alpha)$
metric $\textbf{FDR}_{T}$ and $\textbf{Ineff}_{T}$ $\textbf{Reg}_{T}$
Here, we map the components of online selective generation with partial feedback to those of adversarial bandits to leverage existing algorithms and regret bounds. In particular, each parameter $\tau\in\mathcal{{H}}$ of a selective generator is considered as an arm $\tau\in\mathcal{{H}}$ in adversarial bandits. We learn selective generators from partial feedback $\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq_{E}\mathbf{y}_{t})$, i.e., whether an output of a selective generator $\hat{{S}}$ with a chosen parameter $\tau_{t}$ is incorrect or not, where $\hat{{S}}(\mathbf{x}_{t};\tau_{t})=\texttt{IDK}$ incurs non-zero feedback by the definition. In adversarial bandits, we consider two-step feedback: (1) get feedback from the nature $\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq_{E}\mathbf{y}_{t})$ and (2) use this to compute the loss (6) by the chosen arm $\tau_{t}$. Based on observed loss up to the time horizon $T$, the adversarial bandit learner updates a strategy $p_{t}$ over time to minimize the regret $\textbf{Reg}_{T}$. Then, by leveraging our conversion lemma (Lemma 1), the algorithm provides an FDR bound. See Table 1 for a summary of reduction.
4.2 Algorithm: Exp3 for Online Selective Generation with Partial Feedback Unlocking
As for the adversarial bandit learner, we leverage Exp3 auer2002nonstochastic that is a direct extension of the EW littlestone1994weighted for partial feedback. The key challenge in moving from full to partial feedback is the slow convergence of learning due to the lack of feedback information. To address this, we exploit the unique parameter structure of selective generators to unlock feedback of other arms by a chosen arm, i.e., feedback unlocking. In particular, by the monotonicity of a selection function in $\tau$, i.e., $\hat{{s}}(\mathbf{x})\coloneqq\mathbbm{1}(f_{t}\geq\tau)$ where $f_{t}\coloneqq f(\mathbf{x}_{t},G(\mathbf{x}_{t}))$ and the definition of a selective generator $\hat{{S}}$, the following relation among feedback holds, where $b_{t}\coloneqq\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq_{E}\mathbf{y}_{t})$ is observed feedback by a chosen arm $\tau_{t}$:
[…]
To exploit this feedback unlocking within Exp3, we propose a novel unbiased loss estimator
$\ell_{t}(\tau,\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))\coloneqq\frac{\ell_{t}(\tau,\alpha)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\cdot\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau_{t}}))\in[0,\infty),$ (8)
where $\mathcal{{H}}_{t}({\tau_{t}})\coloneqq\{\tau\in\mathcal{{H}}\mid\tau\leq f_{t}\}$ if $\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}$ and $\mathcal{{H}}_{t}(\tau_{t})\coloneqq\{\tau\in\mathcal{{H}}\mid\tau>f_{t}\}$ otherwise. Note that $\tau\in\mathcal{{H}}_{t}(\tau)$ by its definition and it is an unbiased estimator, i.e., $\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau,\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))=\ell_{t}(\tau,\alpha)$. See (29) in Appendix G.4 for a proof on the unbiasedness and Algorithm 1 for our extension of Exp3 with partial feedback unlocking.
Connection to Exp3. Our algorithm can be reduced to Exp3 by letting $\mathcal{{H}}_{t}({\tau})=\{\tau_{t}\}$ for any $\tau$. Then, we have $\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau_{t}\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})=p_{t}(\tau_{t})$, recovering the Exp3 unbiased estimator in (2) from our unbiased estimator (8).
Connection to EW. Our algorithm can be also reduced to EW. Specifically, let $\mathcal{{H}}_{t}({\tau})=\mathcal{{H}}$ for any $\tau$, i.e., full feedback. Then, we have $\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau_{t}\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})=1$, making our estimator (8) the same as the loss $\ell_{t}$.
Algorithm 1 Exp3 for Online Selective Generation with Partial Feedback Unlocking (ExSUL)
1:procedure ExSUL($T,\mathcal{{H}},\alpha,\lambda,\eta,f,G$)
2:     $w_{1}(\tau)\leftarrow 1/|\mathcal{{H}}|$ for all $\tau\in\mathcal{{H}}$
3:     for $t=1,\ldots,T$ do
4:         Observe $\mathbf{x}_{t}$ or equivalently $f_{t}\coloneqq f(\mathbf{x}_{t},G(\mathbf{x}_{t}))$
5:         Predict $\hat{{S}}(\mathbf{x}_{t};\tau_{t})$ where $\tau_{t}\sim p_{t}(\tau)=\sum_{\tau\in\mathcal{{H}}}\delta(\tau)\cdot{w_{t}(\tau)}/{\sum_{\tau\in\mathcal{{H}}}w_{t}(\tau)}$
[…]
7:         if  $\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}$  then
8:              $\mathcal{{H}}_{t}(\tau_{t})\leftarrow\{\tau\in\mathcal{{H}}\mid\tau\leq f_{t}\}$
9:         else
10:              $\mathcal{{H}}_{t}(\tau_{t})\leftarrow\{\tau\in\mathcal{{H}}\mid\tau>f_{t}\}$
[…]
14:procedure ComputeLoss($\hat{\mathbf{s}},b,\alpha,\lambda$)
15:     $a\leftarrow\mathbbm{1}(\hat{\mathbf{s}}=\texttt{IDK})$
16:     $d\leftarrow\mathbbm{1}(\hat{\mathbf{s}}\neq\texttt{IDK})\cdot b-\alpha\,\mathbbm{1}(\hat{\mathbf{s}}\neq\texttt{IDK})+\alpha$
17:     $\ell\leftarrow a+\lambda d$
18:     return $\ell$
4.3 Regret Bound
Our Algorithm 1 has a sublinear regret bound. From this regret bound, we construct the FDR bound by using Lemma 1. See Appendix G.4 for a proof on our novel regret bound.
Theorem 2.
Let $\ell_{t}(\cdot)\in[0,\ell_{\text{max}}]$. For any $T\in\mathbb{N}$ and finite hypotheses $\mathcal{{H}}$, Algorithm 1 provides the following regret bound if $\eta=\sqrt{{\ln|\mathcal{H}|}\mathbin{/}{\ell_{\text{max}}^{2}T}}$:
$\textbf{Reg}_{T}=\sum_{t=1}^{T}\mathbbm{E}_{\tau,\tau_{t}\sim p_{t}}\ell_{t}(\tau,\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau,\alpha)\leq 2\ell_{\text{max}}\sqrt{T\ln|\mathcal{H}|}$
for any observed loss sequences $\ell_{t}$ generated before learning by an oblivious adversary.
Interestingly, despite of partial feedback, our algorithm achieves the same upper bound as EW with full feedback, i.e., $\ell_{\text{max}}\sqrt{{T\ln|\mathcal{{H}}|}/{2}}$, up to a constant factor due to our unbiased loss estimator (8) with rich information, whereas Exp3 suffers from an additional factor of $\sqrt{|\mathcal{{H}}|}$ (see (3)) due to its unbiased loss estimator with limited information.
4.4 Regret-to-FDR Conversion
We mainly leverage algorithms that minimize the regret $\textbf{Reg}_{T}$ for learning selective generators. Yet, the key requirement for selective generation is the FDR guarantee of a learner at a desired level. To this end, we introduce a novel perspective on the connection of a regret bound to an FDR bound. In particular, any learner, including learners for online learning and bandits, that minimizes the regret (1) controls the FDR (4). This can be achievable as the designed loss (6) is penalized with the regularized parameter $\lambda$ when the FDR loss at each step is higher than $\alpha$ by $d_{t}(\tau,\alpha)$, resulting in controlling the FDR over time $T$ by $\alpha$. The following is a proposed regret-to-FDR conversion lemma. See Appendix G.3 for a proof.
Lemma 1.
Let $T\in\mathbb{N}$, $\alpha\in(0,1)$. If we have a learner that achieves a bounded regret for any loss sequences of (6), then the learner also bounds the FDR, i.e., if $\lambda\geq 1/T^{1/4}$
$\textbf{FDR}_{T}\leq\alpha+\frac{1+\textbf{Reg}_{T}/T}{(1-\textbf{Ineff}_{T})T^{1/4}}.$
Importantly, this lemma is applicable to learning for both full and partial feedback, as it is agnostic to feedback mechanisms. Also, the lemma implies that if $\textbf{Reg}_{T}$ has a sublinear bound, satisfied by most online learners, and $\textbf{Ineff}_{T}$ is constant less than 1, which also mostly holds, then $\textbf{FDR}_{T}$ gets closer to $\alpha$ at most the rate of $\mathcal{{O}}(1/T^{1/4})$, i.e., $\textbf{FDR}_{T}\leq\alpha+\mathcal{{O}}(1/T^{1/4})$. Note that $\textbf{Ineff}_{T}$ does have an effect on the convergence speed by a constant factor. See Section H for discussion on efficiency.
5 Experiments
We empirically justify that ExSUL controls the FDR while maximizing selection efficiency. Moreover, we demonstrate the controllability guarantee holds under diverse environments. In particular, we consider three environments: (1) stochastic, (2) distribution-shift, and (3) interactive environments.
Datasets and Models. We use two datasets for stochastic and distribution-shift environments, $79\mathrm{K}$ Natural Question (NQ) kwiatkowski2019natural and $93\mathrm{K}$ TriviaQA joshi2017triviaqa with two base models, GPT-3.5-turbo openai2023gpt35turbo and LLaMA3.1-8B-Instruct meta2024llama31-8binstruct . We simulate distribution-shift environments by mixing two datasets in diverse ways and an interactive environment by generating interaction between two GPT-3.5-turbo models (see Section E.2 and E.3, respectively for details). To obtain entailment labels and self-consistency scores, we additionally use GPT-3.5-turbo as an entailment model.
Scoring functions. We consider two scoring functions $f_{\texttt{std}}$ and $f_{\texttt{con}}$, introduced in Section 2.1. In short, $f_{\texttt{std}}$ is a standard log-likelihood score defined as the conditional probability of an answer given a question. $f_{\texttt{con}}$ is a self-consistency score of an answer computed via entailment scores (from an entailment model) across samples. We use $f_{\texttt{con}}$ unless specified.
Methods. We consider two baselines, Exp3-SG and No-SG, and one performance upper bound algorithm, EW-SG. In particular, (1) Exp3-SG (Algorithm 5) is an online selective generator with partial feedback by converting Exp3 auer2002nonstochastic for selective generation via our regret-to-FDR conversion lemma. (2) No-SG is a non-selective generator (i.e., $\tau_{t}=0$ always), to serve as a standard use of a generator without abstention. (3) EW-SG (Algorithm 4) is an online selective generator with full feedback by converting this for selective generation via our conversion lemma. As it exploits full feedback, we use it as our empirical performance upper bound.
Metrics. We measure performance of ours along with comparing methods via the FDR, i.e., $\textbf{FDR}_{t}$ and selection inefficiency, i.e., $\textbf{Ineff}_{t}$. Note that No-SG achieves zero selection inefficiency by definition, so we do not explicitly add in figures.
Parameters. A desired FDR level $\alpha$ and time horizon $T$ depend on tasks, but we set $|\mathcal{{H}}|=1\mathrm{K}$ and $\lambda=T^{\nicefrac{{1}}{{4}}}$ by default. Note that $\alpha$ is a desired FDR set by users. This usually depends on tasks, e.g., $\alpha$ is smaller than an error of a base generator on a given task without selection to avoid a trivial FDR.
5.1 Stochastic Environment
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 2: Comparison of selective generation methods under a stochastic environment with LLaMA3.1-8B-Instruct as a generator on TriviaQA ($T=30\mathrm{K},\alpha=0.08$). The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples over $30$ random trials.
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 3: Comparison of selective generation methods under a stochastic environment with GPT-3.5-turbo as a generator on TriviaQA ($T=30\mathrm{K},{\alpha=0.08}$). The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples over $30$ random trials.
In Figure 2 and 3, we observe that ExSUL converges fast under $\alpha$, comparable to EW-SG which has access to full feedback, while Exp3-SG fails to converge under $\alpha$ in most cases. This is because Exp3-SG only observes partial feedback from the chosen arm, requiring a much longer time horizon $T$ for convergence (Figure 23), which supports the regret bound analysis discussed in Section 4.3. These results imply that our partial feedback unlocking is beneficial in a regret bound and faster convergence to the desired FDR. Similar trends are observed in Figure 9 and 11. Additionally, Figure 21 demonstrates that our algorithm effectively controls the FDR under varying $\alpha$ in practice. Note that algorithms with $f_{\texttt{con}}$ have faster convergence than those with $f_{\texttt{std}}$ (e.g., Figure 9 v.s. 11), suggesting that a well-calibrated scoring function can have a positive effect on convergence speed.
5.2 Distribution-shift Environment
In this study, to evaluate the robustness of selective generation we examine three distinct distribution-shift scenarios. First, we create a single instantaneous change by concatenating one large segment from one dataset with an other segment from the other. Second, we induce frequent, periodic shift by interleaving a series of small, fixed-size chunks drawn randomly from each dataset. Third, we provide progressive shift by sampling each example according to a mixing probability that varies linearly. More details of environment setup introduce in Appendix E.2.
In Figure 4,5, and 6, we observes that ExSUL is clearly better in convergence speed than Exp3-SG, while closely follows EW-SG for data distribution shifted environment. Interestingly, in Figure 4, Exp3-SG exhibits a sharp increase in the FDR immediately following the distribution shift, highlighting its sensitivity to sudden changes in data distribution, while the proposed ExSUL does not. In 22, as in the stochastic environments, we also confirmed that our methods find selective generators for a diverse desired FDR $\alpha$, showing theoretical controllability empirically. See additional results on varying generators and shift types in Section F.2 and a longer time horizon in Section 24.
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 4: Comparison of selective generation methods under a distribution-shift environment with GPT-3.5-turbo as a generator ($T=120\mathrm{K},\alpha=0.1$). Here, we consider a single distribution shift from $60\mathrm{K}$-sized TriviaQA to $60\mathrm{K}$-sized NQ, denoted in a dotted vertical line. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 5: Comparison of selective generation methods under a distribution-shift environment with GPT-3.5-turbo as a generator ($T=120\mathrm{K},\alpha=0.1$). We consider multiple distribution shifts by alternating $10\mathrm{K}$-sized TriviaQA and NQ 6 times (i.e., TriviaQA, NQ, TriviaQA, NQ and so forth), where shifting points are visualized in dotted vertical lines. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 6: Comparison of selective generation methods under a distribution-shift environment with GPT-3.5-turbo as a generator ($T=60\mathrm{K},\alpha=0.1$). We consider a distribution shift by gradually shifting from TriviaQA to NQ over time. The violin plots are drawn with randomly chosen $60\mathrm{K}$ samples with $30$ random trials.
5.3 Interactive Environment
We empirically demonstrate the efficacy of our proposed ExSUL in real-world interactive applications. In particular, we simulate an interactive environment by implementing a user-acting agent, a question-answering agent, and an evaluating agent by states. These agents interact over multiple turns, forming a dialog-like sequence of states that mimics realistic user interactions, and our selective generation method ExSUL decides whether to abstain from generation as shown in Figure 1. In Figure 7, we observe that ExSUL consistently controls the FDR under $\alpha$, despite the instability of $\textbf{FDR}_{t}$ due to the shifting dynamics of question-answering. This observation demonstrates that our method ExSUL is robust under frequently shifting distributions over time. In Figure 1, ExSUL successfully abstains from incorrect answers, helping to achieve the desired FDR level $\alpha$.
(a) FDR over time
(b) Inefficiency over time
Figure 7: $\textbf{FDR}_{t}$ and $\textbf{Ineff}_{t}$ for ExSUL under an interactive environment with GPT-3.5-Turbo as a question-answering agent ($T=15{,}000$, $\alpha=0.25$). A user agent continuously creates questions based on changing context, which simulate distribution shifts.
6 Conclusion
We propose a learning algorithm for online selective generation with partial feedback, which provides an FDR controllability guarantee, while maximizing selection efficiency. To this end, we leverage the well-developed adversarial bandit problem. In particular, we reduce selective generation to adversarial bandits and extend the Exp3 algorithm for adversarial bandits by leveraging a technique called partial feedback unlocking, which addresses the lack of feedback information due to the nature of partial feedback. Finally, we introduce a novel Regret-to-FDR conversion lemma such that we can use any regret minimization algorithm, including our extended Exp3 for selective generation to control an FDR at a desired level. We theoretically and empirically justify our method by providing an efficient regret bound and evaluating diverse learning environments, including stochastic, distribution-shift, and interactive environments.
Limitations. As for the main limitation of our method, we assume that an oblivious adversary generates data sequences, i.e., a data sequence is drawn before learning, so our FDR guarantee may not hold if an stronger adversary generates data sequences adaptively.
Acknowledgment
We appreciate constructive feedback by anonymous reviewers. This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH) (5%); RS-2024-00457882, National AI Research Lab Project (20%); No. RS-2024-00509258 and No. RS-2024-00469482, Global AI Frontier Lab (40%); RS-2025-00560062 (35%)).
Appendix A Additional Related Work
Constrained Online Learning Several works have studied online learning with constraints. [19] first proposes long-term constraints in online convex optimization with time-invariant constraints, which allows to have the constraint function $g_{t}$ to be violated at each time step, i.e., $g(\mathbf{x}_{t})>0$, but eventually satisfied, i.e., $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$. Some other works solve a similar problem but with time-varying constraints, which is more challenging. [20, 21] solve a network resource allocation task by viewing this as an constrained optimization problem. Especially, [21] solves the problem without a convex assumption on loss. [22] gives analysis on optimal algorithms for online convex optimization with adversarial constraints. We similarly consider a FDR guarantee to be satisfied over time but with partial feedback.
Conformal Prediction with Bandit Feedback There is little work about selective or conformal prediction in bandit settings. [23] proposes an online conformal prediction method but with semi-bandit feedback, where we can observe the true label only if it is predicted. Our method follows the practical motivation in having partial feedback idea but we consider selective generation.
Appendix B Additional Preliminary
B.1 Online Learning and EW Algorithm
Online learning considers to design a learner that adapts to any sequence of full feedback on hypotheses without any or with little assumption on data generation process. In particular, the goal of an online learner is to find a distribution $p_{t}$ over hypotheses with full feedback that minimizes the regret $\textbf{Reg}_{T}$.
Along with many online learning algorithms, we introduce Exponential Weighting (EW) [14], which minimizes the regret with finite hypotheses (Algorithm 2). In particular, EW maintains the goodness of weights for each hypothesis in terms of cumulative loss and updates them from feedback. The feedback of a hypothesis $\tau_{t}$ at time $t$ is represented by loss $\ell_{t}(\tau_{t})$ but we can compute this loss for all hypotheses other than a chosen hypothesis $\tau_{t}$ as we directly observe a true label $\mathbf{y}_{t}$ at the same time $t$. This EW algorithm satisfies the following regret bound. See Appendix G.1 for a proof.
Theorem 3.
[14, 13] Let $\ell_{t}(\cdot)\in[0,\ell_{\text{max}}]$. For any $T\in\mathbb{N}$ and finite hypotheses $\mathcal{{H}}$, Algorithm 2 provides the following regret bound if $\eta=\sqrt{{8\ln|\mathcal{{H}}|}/{(\ell_{\text{max}}^{2}T)}}$:
$\textbf{Reg}_{T}\leq\ell_{\text{max}}\sqrt{{T\ln|\mathcal{{H}}|}/{2}}=\mathcal{{O}}(\ell_{\text{max}}\sqrt{T\ln|\mathcal{{H}}|}).$ (9)
We will leverage and extend this EW for learning selective generators with full feedback.
Appendix C Additional Method
C.1 Online Selective Generation with Full Feedback
Here, we specifically leverage the conventional exponential weighting (EW) algorithm [14] for online learning under adaptive adversaries to re-purpose it for addressing online selective generation with full feedback. This is an oracle method with the performance upper bound of our method with partial feedback. To this end, we provide the reduction of online selective generation with full feedback to the online learning problem. Then, we introduce a modified EW algorithm for selective generation, called EW-SG, followed by its regret bound. Finally, we leverage the proposed conversion lemma, which converts the regret bound of any online learning algorithms into the FDR bound, to show the FDR controllability of EW-SG at a desired level.
C.1.1 Reduction: From Online Selective Generation to Online Learning
Table 2: From online selective generation to online learning
online selective generation online learning
models selective generators $\mathcal{{H}}$ hypotheses $\mathcal{{H}}$
feedback $\mathbbm{1}(G(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t})$ $\ell_{t}(\tau,\alpha)$
metric $\textbf{FDR}_{T}$ and $\textbf{Ineff}_{T}$ $\textbf{Reg}_{T}$
We convert online selective generation with full feedback to conventional online learning (See Table 2 for a summary). In selective generation, models are defined as a set of selective generators $\hat{{S}}$ or equivalently a set of selection parameters $\mathcal{{H}}$, considered as hypotheses $\mathcal{{H}}$ in online learning. This model is sequentially updated by leveraging online feedback. In particular, we consider feedback whether a generated answer $G(\mathbf{x}_{t})$ is semantically incorrect with respect to a true answer $\mathbf{y}_{t}$, i.e., $\mathbbm{1}(G(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t})$. This essentially provides full feedback for any $\tau\in\mathcal{{H}}$ as it is the same as $\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau)\neq_{E}\mathbf{y}_{t})$ by the definition of $\hat{{S}}$ if $\hat{{S}}(\mathbf{x}_{t};\tau)\neq\texttt{IDK}$. This feedback is converted into loss in online learning, as defined in (6). Note that $d_{t}$ in the loss leverages the feedback $\mathbbm{1}(G(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t})$ by the definition. Finally, our goal of finding a learning algorithm, which controls the FDR by $\alpha$ and minimizes selection inefficiency in online selective generation, is reduced to finding an online learning algorithm that minimizes the regret $\textbf{Reg}_{T}$ with any loss sequences $\ell_{t}(\tau,\alpha)$ in (6).
C.1.2 Algorithm and Its Regret Bound
We re-purpose EW for selective generation called EW-SG. This mainly uses the special loss $\ell_{t}(\tau,\alpha)$ in (6) for selective generation. See Algorithm 4 for details. The corresponding regret bound is directly obtained from Theorem 3 where $\ell_{\text{max}}\coloneqq\max(\lambda,1+\lambda\alpha)\leq 1+\lambda$.
C.2 Regret-to-FDR Conversion
Once we devise a regret minimization algorithm, i.e., EW-SG, the regret guarantee is simply interpreted as the FDR guarantee by Lemma 1. This eventually provides the FDR bound for EW-SG.
Appendix D Algorithms
Algorithm 2 Exponential Weighting (EW) [14, 13]
1:procedure EW($T,\mathcal{{H}},\eta$)
2:     $w_{1}\leftarrow(1/|\mathcal{{H}}|,\dots,1/|\mathcal{{H}}|)$
3:     for $t=1,\ldots,T$ do
4:         Observe $\mathbf{x}_{t}$
5:         Predict $\hat{\mathbf{y}}(\mathbf{x}_{t};\tau_{t})$ for $\tau_{t}\sim p_{t}(\tau)=\sum_{\tau\in\mathcal{{H}}}\delta(\tau)\cdot w_{t}(\tau)/\sum_{\tau\in\mathcal{{H}}}w_{t}(\tau)$
6:         Observe $\mathbf{y}_{t}$
7:         Update $w_{t+1}(\tau)\propto w_{t}(\tau)\exp(-\eta\ell(\mathbf{y}_{t},\hat{\mathbf{y}}(\mathbf{x}_{t};\tau)))$ for all $\tau\in\mathcal{{H}}$
Algorithm 3 Exponential-weight algorithm for Exploration and Exploitation (Exp3) [7]
1:procedure Exp3($T,\mathcal{{H}},\eta$)
2:     $w_{1}\leftarrow(1/|\mathcal{{H}}|,\dots,1/|\mathcal{{H}}|)$
3:     for $t=1,\ldots,T$ do
4:         Choose an arm $\tau_{t}\sim p_{t}(\tau)=\sum_{\tau\in\mathcal{{H}}}\delta(\tau)\cdot w_{t}(\tau)/\sum_{\tau\in\mathcal{{H}}}w_{t}(\tau)$
[…]
7:         Update $w_{t+1}(\tau)\propto w_{t}(\tau)\cdot\exp(-\eta\hat{\ell}_{t}(\tau))$ for all $\tau\in\mathcal{{H}}$
Algorithm 4 Exponential Weighting for Online Selective Generation (EW-SG)
1:procedure EW-SG($T,\mathcal{{H}},\alpha,\lambda,\eta$)
2:     $w_{1}(\tau)\leftarrow 1/|\mathcal{{H}}|$ for all $\tau\in\mathcal{{H}}$
3:     for $t=1,\ldots,T$ do
4:         Observe $\mathbf{x}_{t}$
5:         Predict $\hat{{S}}_{t}(\mathbf{x}_{t};\tau_{t})$ where $\tau_{t}\sim p_{t}=\sum_{\tau\in\mathcal{{H}}}w_{t}(\tau)\cdot\delta(\tau)$
6:         Observe $\mathbf{y}_{t}$ thus observe $\mathbbm{1}\left(G(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t}\right)$
7:         $a_{t}(\tau)\leftarrow\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau)=\texttt{IDK})$
8:         $d_{t}(\tau,\alpha)\leftarrow\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau)\neq\texttt{IDK}\wedge G(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t})-\alpha\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau)\neq\texttt{IDK}))+\alpha$
9:         $\ell_{t}\left(\tau,\alpha\right)\leftarrow a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)$ for all $\tau\in\mathcal{{H}}$
10:         Update $w_{t+1}(\tau)\propto w_{t}(\tau)\exp\left\{-\eta\cdot\ell_{t}\left(\tau,\alpha\right)\right\}$ for all $\tau\in\mathcal{{H}}$
Algorithm 5 Exp3 for Online Selective Generation with Partial Feedback (Exp3-SG)
1:procedure Exp3-SG($T,\mathcal{{H}},\alpha,\lambda,\eta$)
2:     $w_{1}(\tau)\leftarrow 1/|\mathcal{{H}}|$ for all $\tau\in\mathcal{{H}}$
3:     for $t=1,\ldots,T$ do
4:         Observe $\mathbf{x}_{t}$
5:         Predict $\hat{{S}}_{t}(\mathbf{x}_{t};\tau_{t})$ where $\tau_{t}\sim p_{t}(\tau)=\sum_{\tau\in\mathcal{{H}}}\frac{w_{t}(\tau)}{\sum_{\tau\in\mathcal{{H}}}w_{t}(\tau)}\cdot\delta(\tau)$ ()arm $\left(\triangleright\right)\text{arm selection}$
[…]
10:procedure ComputeLoss($\hat{\mathbf{s}},b,\alpha,\lambda$)
11:     $a\leftarrow\mathbbm{1}(\hat{\mathbf{s}}=\texttt{IDK})$
12:     $d\leftarrow\mathbbm{1}(\hat{\mathbf{s}}\neq\texttt{IDK})\cdot b-\alpha\,\mathbbm{1}(\hat{\mathbf{s}}\neq\texttt{IDK})+\alpha$
13:     $\ell\leftarrow a+\lambda d$
14:     return $\ell$
Appendix E Experiment Setup
E.1 Computing Environment
We use 4 NVIDIA A100 80GB with 128 CPUs for experiments.
E.2 Distribution-shift Environment Setup
We describe the experimental setup used to evaluate the robustness of selective generation by three distinct distribution-shift scenarios. We detail the chunking strategy and the procedure for constructing each shift environment in the following.
[…]
Gradual Shift. This setup generates a smooth distribution change over time by drawing individual samples probabilistically between NQ and TriviaQA, with the sampling probability gradually shifting from pure NQ (or TriviaQA) at the start to pure TriviaQA (or NQ) at the end. In particular, given a multinomial distribution $[\frac{t}{T},T-\frac{t}{T}]$ for each dataset that changes over time, we first randomly choose a dataset which follows this multinomial distribution, and then we sample one question-answering pair in the sampled dataset. Over the full sequence length, this creates linear interpolation between the two datasets, enabling analysis of model performance under a progressive distribution shift.
E.3 Interactive Environment Setup
For the interactive environment, we consider dialog-based conversations. In particular, GPT-3.5-Turbo serves as a question-answering agent, while we utilize two GPT-4o APIs, where one is a user-acting agent to generate questions, and another is an evaluating agent, having assess to the responses.
To instruct the user-acting agent to generate questions, we use the SQuAD [24] dataset, which consists of many question-answer pairs, each associated with a gold context. Here, we discard the QA pairs and only use $20,962$ context passages, which vary over a range of topics (e.g., science, history, celebrities, entertainment, etc.), forming a diverse set of natural distributions. During the simulation, we randomly select a context from the dataset and provide it to the user-acting agent, which is prompted to continuously generate questions based on the given context. (e.g., Figure 1). To mimic a dynamically shifting environment, e.g., the user’s interest shifts over time, we further instruct the user-acting agent to change the context during the interaction when it determines it is done with the current one (e.g., "When you are done with the current context, say ’SHIFT’ to request a new context.") or after at most 10 question-answering turns. To design this custom workflow, we implement the core functions of the simulation, including memory handling and multi-turn conversation, using LangGraph, a framework built on LangChain [25].
Appendix F Additional Experiments
F.1 Stochastic Environment
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 8: Comparison of selective generation methods under a stochastic environment with GPT-3.5-turbo as a generator on NQ ($T=30\mathrm{K},{\alpha=0.25}$). The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples over $30$ random trials.
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 9: Comparison of selective generation methods under a stochastic environment with LLaMA3.1-8B-Instruct as a generator on NQ ($T=30\mathrm{K},{\alpha=0.25}$). The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples over $30$ random trials. Comparison of selective generation methods under stochastic environments with varying feedback.
(a) FDR over time
(b) Inefficiency over time
Figure 10: Comparison of selective generation methods under a stochastic environment with GPT-3.5-turbo as a generator on NQ ($T=30\mathrm{K},{\alpha=0.25}$) and $f_{\texttt{std}}$ as our scoring function. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples over $30$ random trials.
(a) FDR over time
(b) Inefficiency over time
Figure 11: Comparison of selective generation methods under a stochastic environment with LLaMA3.1-8B-Instruct as a generator on NQ ($T=30\mathrm{K},{\alpha=0.25}$) and $f_{\texttt{std}}$ as our scoring function. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples over $30$ random trials.
F.2 Distribution-shift Environment
F.2.1 Single Shift
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 12: Comparison of selective generation methods under a distribution-shift environment with GPT-3.5-turbo as a generator ($T=120\mathrm{K},\alpha=0.1$). Here, we consider a single distribution shift from $60\mathrm{K}$-sized NQ to $60\mathrm{K}$-sized TriviaQA, denoted in a dotted vertical line. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 13: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=120\mathrm{K},\alpha=0.1$). Here, we consider a single distribution shift from $60\mathrm{K}$-sized NQ to $60\mathrm{K}$-sized TriviaQA, denoted in a dotted vertical line. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 14: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=120\mathrm{K},\alpha=0.1$). Here, we consider a single distribution shift from $60\mathrm{K}$-sized TriviaQA to $60\mathrm{K}$-sized NQ, denoted in a dotted vertical line. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
F.2.2 Alternating Shift
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 15: Comparison of selective generation methods under a distribution-shift environment with GPT-3.5-turbo as a generator ($T=120\mathrm{K},\alpha=0.1$). We consider multiple distribution shifts by alternating $10\mathrm{K}$-sized NQ and TriviaQA 6 times (i.e., NQ, TriviaQA, NQ, TriviaQA and so forth), where shifting points are visualized in dotted vertical lines. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 16: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=120\mathrm{K},\alpha=0.1$). We consider multiple distribution shifts by alternating $10\mathrm{K}$-sized NQ and TriviaQA 6 times (i.e., NQ, TriviaQA, NQ, TriviaQA and so forth), where shifting points are visualized in dotted vertical lines. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 17: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=120\mathrm{K},\alpha=0.1$). We consider multiple distribution shifts by alternating $10\mathrm{K}$-sized TriviaQA and NQ 6 times (i.e., TriviaQA, NQ, TriviaQA, NQ and so forth), where shifting points are visualized in dotted vertical lines. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
F.2.3 Gradual Shift
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 18: Comparison of selective generation methods under a distribution-shift environment with GPT-3.5-turbo as a generator ($T=60\mathrm{K},\alpha=0.1$). We consider a distribution shift by gradually shifting from NQ to TriviaQA over time. The violin plots are drawn with randomly chosen $60\mathrm{K}$ samples with $30$ random trials.
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 19: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=60\mathrm{K},\alpha=0.1$). We consider a distribution shift by gradually shifting from NQ to TriviaQA over time. The violin plots are drawn with randomly chosen $60\mathrm{K}$ samples with $30$ random trials.
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 20: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=60\mathrm{K},\alpha=0.1$). We consider a distribution shift by gradually shifting from TriviaQA to NQ over time. The violin plots are drawn with randomly chosen $60\mathrm{K}$ samples with $30$ random trials.
F.3 Analysis on Time Horizon and Desired FDR Parameters
(a) NQ dataset along with GPT-3.5-Turbo
(b) NQ dataset along with LLaMA3.1-8B
(c) TriviaQA dataset along with GPT-3.5-Turbo
(d) TriviaQA dataset along with LLaMA3.1-8B
Figure 21: The violin plots for the ablation study over varying $\alpha$ on datasets under stochastic environments. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples over 30 random trials. Here, converged FDRs mostly go along with target risk $\alpha$ (i.e., violin plots are below of the diagonal line), which empirically demonstrates that our algorithm finds efficient selective generators with FDR guarantees. Note that $\textbf{Err}_{T}$ refers to the error rate on a given dataset, i.e., $\textbf{Err}_{T}\coloneqq\sum_{t=1}^{T}\mathbbm{1}\left(G(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t}\right)/T$ and the FDRs of selective generators should be lower than $\textbf{Err}_{T}$ to be meaningful in controlling the FDRs.
(a) TrviaQA shifts to NQ with GPT-3.5-Turbo
(b) TrviaQA shifts to NQ with LLaMA3.1-8B
(c) NQ shifts to TriviaQA with GPT-3.5-Turbo
(d) NQ shifts to TriviaQA with LLaMA3.1-8B
Figure 22: The violin plots for the ablation study over varying $\alpha$ on datasets under distribution-shift environments. The violin plots are drawn with randomly chosen dataset with single distribution shift between $60\mathrm{K}$-sized TriviaQA and $60\mathrm{K}$-sized NQ, over 30 random trials. Here, converged FDRs go along with target risk $\alpha$ (i.e., violin plots are below of the diagonal line), which empirically demonstrates the our algorithm finds an efficient selective generators with FDR guarantees. Note that $\textbf{Err}_{T}$ refers to the error rate on a given dataset, i.e., $\textbf{Err}_{T}\coloneqq\sum_{t=1}^{T}\mathbbm{1}\left(G(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t}\right)/T$ and the FDRs of selective generators should be lower than $\textbf{Err}_{T}$ to be meaningful in controlling the FDRs.
(a) NQ dataset along with GPT-3.5-Turbo
(b) NQ dataset along with LLaMA3.1-8B
(c) TriviaQA dataset along with GPT-3.5-Turbo
(d) TriviaQA dataset along with LLaMA3.1-8B
Figure 23: Additional $\textbf{FDR}_{t}$ comparisons of selective generation methods under stochastic environments with a longer time horizon. Here, we repeat the datasets to evaluate whether Exp3-SG converges over a long horizon ($T=1,500\mathrm{K}$). As shown in the plots, Exp3-SG exhibits bare convergence after a prolonged horizon $T$.
Figure 24: Additional $\textbf{FDR}_{t}$ comparison of selective generation methods under distribution-shift environments with a longer time horizon. We consider a single distribution shift from $2000\mathrm{K}$-sized TriviaQA to $2000\mathrm{K}$-sized NQ, denoted in a dotted vertical line, along with GPT-3.5-turbo as our base generator ($T=4000\mathrm{K},\alpha=0.1$). As can be seen, Exp3-SG eventually satisfies a FDR even under the shift.
F.4 Ablation Study
(a) FDR over time
(b) Inefficiency over time
Figure 25: Comparison of ExSUL and ExSUL without the inefficiency loss $a_{t}$ (6) for the ablation study, with GPT-3.5-turbo as a generator on TriviaQA ($T=30\mathrm{K},\alpha=0.25$). Removing $a_{t}$ in ExSUL clearly leads to convergence to more inefficient solution. This behavior contrasts with the faster convergence of EW-SG compared to ExSUL or Exp3-SG , e.g., EW-SG and ExSUL eventually converges to similar levels of inefficiency in Figure 3. However, as seen in this plot, ExSUL without $a_{t}$ visibly converges to a much more inefficient generator, which demonstrates our algorithm provides an FDR guarantee while maximizing selection efficiency. Note that $\ell_{\text{max}}$ is set to $\lambda$, and the learning rate is also adjusted accordingly.
Appendix G Proofs
G.1 A Proof of Theorem 3
We follow the standard proof [13], followed by extending it for a loss range with $[0,\ell_{\text{max}}]$.
Let the cumulative loss of the $i$-th hypothesis $\tau_{i}\in\mathcal{{H}}$ be $L_{t}^{i}\coloneqq\sum_{s=1}^{t}\ell_{t}(\tau_{i})$ and $W_{t}\coloneqq\sum_{i=1}^{|\mathcal{{H}}|}\exp\left\{-\eta L_{t}^{i}\right\}$.
The upper bound of $\ln\frac{W_{t}}{W_{t-1}}$. We have the following upper bound of $\ln\frac{W_{t}}{W_{t-1}}$:
$\ln\frac{W_{t}}{W_{t-1}}$ $=\ln\frac{\sum_{i=1}^{|\mathcal{{H}}|}\exp\left\{-\eta L_{t}^{i}\right\}}{\sum_{i=1}^{|\mathcal{{H}}|}\exp\{-\eta L_{t-1}^{i}\}}$
$=\ln\sum_{i=1}^{|\mathcal{{H}}|}\exp\left\{-\eta\ell_{t}(\tau_{i})\right\}\frac{\exp\{-\eta L_{t-1}^{i}\}}{\sum_{i=1}^{|\mathcal{{H}}|}\exp\{-\eta L_{t-1}^{i}\}}$
$=\ln\mathbbm{E}_{\tau_{t}\sim p_{t}}\exp\left\{-\eta\ell_{t}(\tau_{t})\right\}$
$\leq-\eta\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t})+\frac{\eta^{2}\ell_{\text{max}}^{2}}{8},$ (10)
where (10) holds by the Hoeffding’s lemma and the fact that $\ell_{t}(\tau_{i})\in[0,\ell_{\text{max}}]$. Thus, we have
$\ln\frac{W_{T}}{W_{0}}=\sum_{t=1}^{T}\ln\frac{W_{t}}{W_{t-1}}\leq-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t})+\frac{\eta^{2}\ell_{\text{max}}^{2}T}{8}.$
The lower bound of $\ln\frac{W_{t}}{W_{t-1}}$. We have the lower bound of $\ln\frac{W_{t}}{W_{t-1}}$ as follows:
$\ln{\frac{W_{T}}{W_{0}}}=\ln\sum_{i=1}^{|\mathcal{{H}}|}\exp\{-\eta L_{T}^{i}\}-\ln|\mathcal{{H}}|$ $\geq\ln\left(\max_{i\in\{1,\dots,|\mathcal{{H}}|\}}\exp\{-\eta L^{i}_{T}\}\right)-\ln|\mathcal{{H}}|$
$=-\eta\min_{i\in\{1,\dots,|\mathcal{{H}}|\}}L^{i}_{T}-\ln|\mathcal{{H}}|$
$=-\eta\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau)-\ln|\mathcal{{H}}|.$
Combine the lower and upper bounds. If we combine the lower bound and the upper bound, we obtain the following inequality:
$\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell(\tau_{t})-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau)\leq\frac{\eta\ell_{\text{max}}^{2}T}{8}+\frac{\ln|\mathcal{{H}}|}{\eta}.$
Since the above upper bound of the regret is convex in $\eta$, we can obtain the optimal value of $\eta=\frac{2}{\ell_{\text{max}}}\sqrt{\frac{2\ln|\mathcal{{H}}|}{T}}$ that minimizes the upper bound, which completes the proof.
G.2 A Proof of Theorem 1
This proof follows the standard proof [13] but with a non-trivial loss range $[0,\ell_{\text{max}}]$.
Recall that $\ell_{t}(\tau)\in[0,\ell_{\text{max}}]$ and its unbiased estimator ~$\tilde{\ell}_{t}(\tau\mid\tau_{t})\in[0,\infty)$, as defined in (2). Additionally, we define
$W_{t}\coloneqq\sum_{i=1}^{|\mathcal{{H}}|}\text{exp}(-\eta\tilde{L}_{t}^{i})\quad\text{where}\quad\tilde{\ell}_{t}^{i}\coloneqq\tilde{\ell}_{t}(\tau_{i}\mid\tau_{t})\text{~{}and~{}}\tilde{L}_{t}^{i}\coloneqq\sum_{s=1}^{t}\tilde{\ell}_{s}^{i}.$
The upper bound of potentials.
We find an upper bound of $\mathbbm{E}\ln\frac{W_{T}}{W_{0}}$. To this end, we first consider an upper bound of $\mathbbm{E}\ln\frac{W_{t}}{W_{t-1}}$.
$\ln\frac{W_{t}}{W_{t-1}}$ $=\ln\frac{\sum_{i=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t}^{i})}{\sum_{i=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t-1}^{i})}$
$=\ln\frac{\sum_{i=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t-1}^{i})\,\exp(-\eta\tilde{\ell}_{t}^{i})}{\sum_{i=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t-1}^{i})}$
$=\ln\sum_{i=1}^{|\mathcal{H}|}\frac{\exp(-\eta\tilde{L}_{t-1}^{i})}{\sum_{j=1}^{|\mathcal{H}|}\exp(-\eta\tilde{L}_{t-1}^{j})}\exp(-\eta\tilde{\ell}_{t}^{i})$
$=\ln{\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})}\exp(-\eta\tilde{\ell}_{t}^{i})$ (11)
$\leq\ln\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\left(1-\eta\tilde{\ell}_{t}^{i}+\frac{\eta^{2}(\tilde{\ell}_{t}^{i})^{2}}{2}\right)$ (12)
$=\ln\left(1-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})(\tilde{\ell}_{t}^{i})^{2}\right)$
$\leq-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})(\tilde{\ell}_{t}^{i})^{2}$ (13)
$=-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\left(\frac{\ell_{t}^{i}}{p_{t}(\tau_{i})}\cdot\mathbbm{1}(\tau_{t}=\tau_{i})\right)\tilde{\ell}_{t}^{i}$ (14)
$=-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\ell_{t}^{t}\tilde{\ell}_{t}^{t}$
$=-\eta\mathbbm{E}_{\tau_{i}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\ell_{t}^{t}\tilde{\ell}_{t}^{t}.$ (15)
Here, (11) follows directly from the definition of the sampling distribution $p_{t}(\tau)$. (12) is obtained by applying the second-order Taylor bound $e^{x}\leq 1+x+{x^{2}}/{2}$ which is valid for $x\leq 0$, together with the observation that $-\eta\tilde{\ell}_{t}^{i}\leq 0$. (13) holds by the standard logarithmic bound $\ln(1+x)\leq x$ for any $x\in\mathbb{R}$. Finally, (14) follows from the definition of our estimator.
Then, we have the following:
$\mathbbm{E}\left[\ln\frac{W_{T}}{W_{0}}\right]$ $=\mathbbm{E}\left[\sum_{t=1}^{T}\ln\frac{W_{t}}{W_{t-1}}\right]$
$\leq-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}^{t}\tilde{\ell}_{t}^{t}$ (16)
$=-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{t=1}^{T}\sum_{\tau_{t}\in\mathcal{{H}}}p_{t}(\tau_{t})\ell_{t}(\tau_{t})\left(\frac{\ell_{t}(\tau_{t})}{p_{t}(\tau_{t})}\mathbbm{1}(\tau_{t}=\tau_{t})\right)$
$=-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{t=1}^{T}\sum_{\tau\in\mathcal{{H}}}p_{t}(\tau)\ell_{t}(\tau)\frac{\ell_{t}(\tau)}{p_{t}(\tau)}$
$=-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{t=1}^{T}\sum_{\tau\in\mathcal{{H}}}\ell_{t}(\tau)^{2}$
$\leq-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\ell_{\text{max}}^{2}T|\mathcal{{H}}|$ (17)
After the telescoping sum, (16) applies the exponential-weights bound obtained in the previous step to each factor $\ln\frac{W_{t}}{W_{t-1}}$. Finally, (17) holds as $\ell_{t}(\cdot)$ is bounded by $\ell_{\text{max}}$.
The lower bound of potentials.
Then, we find a lower bound of $\mathbbm{E}\ln\frac{W_{T}}{W_{0}}$. In particular, we have
$\mathbbm{E}\left[\ln{\frac{W_{T}}{W_{0}}}\right]$ $=\mathbbm{E}\left[\ln\sum_{i=1}^{|\mathcal{{H}}|}\exp\{-\eta\tilde{L}_{T}^{i}\}\right]-\ln|\mathcal{{H}}|$
$\geq\mathbbm{E}\left[\ln\left(\max_{i\in\{1,\dots,|\mathcal{{H}}|\}}\exp\{-\eta\tilde{L}^{i}_{T}\}\right)\right]-\ln|\mathcal{{H}}|$
$=\mathbbm{E}\left[-\eta\min_{i\in\{1,\dots,|\mathcal{{H}}|\}}\tilde{L}^{i}_{T}\right]-\ln|\mathcal{{H}}|$
$=\mathbbm{E}\left[-\eta\min_{i\in\{1,\dots,|\mathcal{{H}}|\}}\sum_{t=1}^{T}\tilde{\ell}^{i}_{t}\right]-\ln|\mathcal{{H}}|$
$=\mathbbm{E}\left[-\eta\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\tilde{\ell}_{t}(\tau|\tau_{t})\right]-\ln|\mathcal{{H}}|$
$\geq-\eta\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\tilde{\ell}_{t}(\tau|\tau_{t})-\ln|\mathcal{{H}}|$
$=-\eta\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau)-\ln|\mathcal{{H}}|.$
Combine the lower and upper bounds.
Finally, we combine the lower and upper bound of $\mathbbm{E}\ln\frac{W_{T}}{W_{0}}$ to derive our regret bound. In particular, we have
$\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}-\min_{i\in\{1,\dots,|\mathcal{{H}}|\}}\sum_{t=1}^{T}\ell_{t}(\tau_{i})\leq\frac{\eta}{2}\ell_{\text{max}}^{2}T|\mathcal{{H}}|+\frac{\ln|\mathcal{{H}}|}{\eta}.$
Substituting the optimal learning rate $\eta=\sqrt{\frac{2\ln|\mathcal{H}|}{\ell_{\max}^{2}T|\mathcal{H}|}}$ into the preceding bound yields the claimed regret guarantee, which completes the proof.
G.3 A Proof of Lemma 1
We derive a lower bound of $\textbf{Reg}_{T}$ which consists of $\textbf{FDR}_{T}$. In particular, from the definition of $\textbf{Reg}_{T}$ and loss $\ell_{t}(\tau_{,}\alpha)$ in (6), we have the following:
$\textbf{Reg}_{T}$ $=\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\Big{[}a_{t}(\tau_{t})+\lambda d_{t}(\tau_{t},\alpha)\Big{]}-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\left[a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)\right]$
$\geq\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\Big{[}a_{t}(\tau_{t})+\lambda d_{t}(\tau_{t},\alpha)\Big{]}-\sum_{t=1}^{T}a_{t}(\overline{\tau})-\sum_{t=1}^{T}\lambda d_{t}(\overline{\tau},\alpha)$ (18)
$=\sum_{t=1}^{T}\left[\mathbbm{E}_{\tau_{t}\sim p_{t}}a_{t}(\tau_{t})-a_{t}(\overline{\tau})\right]+\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\lambda d_{t}(\tau_{t},\alpha)-\sum_{t=1}^{T}\lambda d_{t}(\overline{\tau},\alpha)$
$\geq-T+\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\lambda d_{t}(\tau_{t},\alpha)-\sum_{t=1}^{T}\lambda d_{t}(\overline{\tau},\alpha),$ (19)
where (18) holds as $\overline{\tau}=\arg\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\lambda d_{t}({\tau},\alpha)$, implying
$\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\left[a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)\right]\leq\sum_{t=1}^{T}\left[a_{t}(\overline{\tau})+\lambda d_{t}(\overline{\tau},\alpha)\right]$
and (19) holds as $a_{t}(\tau)\in[0,1]$ for any $t$ and $\tau$.
This implies that
$\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}d_{t}(\tau_{t},\alpha)$ $\leq\frac{T+\textbf{Reg}_{T}}{\lambda}+\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}d_{t}({\tau},\alpha)\leq\frac{T+\textbf{Reg}_{T}}{T^{\nicefrac{{1}}{{4}}}}+\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}d_{t}({\tau},\alpha)$
if $\lambda\geq T^{\nicefrac{{1}}{{4}}}$. Then, by the definition of $d_{t}(\tau_{t},\alpha)$, we have
$\frac{1}{T}\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\left[\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq_{E}\mathbf{y}_{t}\right)-\alpha\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\right)+\alpha\right]$
$\leq\frac{1+\textbf{Reg}_{T}/T}{T^{\nicefrac{{1}}{{4}}}}+\min_{\tau\in\mathcal{{H}}}\frac{1}{T}\sum_{t=1}^{T}d_{t}({\tau},\alpha),$
which implies
$\frac{1}{T}\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}$ $\left[\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq_{E}\mathbf{y}_{t}\right)-\alpha\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\right)\right]$
$\leq\frac{1+\textbf{Reg}_{T}/T}{T^{\frac{1}{4}}}+\min_{\tau\in\mathcal{{H}}}\frac{1}{T}\sum_{t=1}^{T}d_{t}({\tau},\alpha)-\alpha$
$\leq\frac{1+\textbf{Reg}_{T}/T}{T^{\frac{1}{4}}}$ (20)
where (20) holds as $\tau^{\prime}=1$ (i.e., always abstaining) satisfies the following:
$\min_{\tau\in\mathcal{{H}}}\frac{1}{T}\sum_{t=1}^{T}d_{t}({\tau},\alpha)\leq\frac{1}{T}\sum_{t=1}^{T}d_{t}({\tau^{\prime}},\alpha)=\alpha.$
Thus, by dividing $1-\textbf{Ineff}_{T}=\frac{1}{T}\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK})$ for both sides of (20), we have
$\textbf{FDR}_{T}=\frac{\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq_{E}\mathbf{y}_{t}\right)}{\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\right)}\leq\alpha+\frac{1+\textbf{Reg}_{T}/T}{(1-\textbf{Ineff}_{T})T^{\nicefrac{{1}}{{4}}}}$
which completes the proof.
G.4 A Proof of Theorem 2
We derive a novel regret bound of our algorithm. The proof technique follows that of the Exp3 regret bound except that we use a novel unbiased estimator for feedback unlocking.
First, we introduce some properties of $\mathcal{{H}}_{t}(\tau_{t})$ in our algorithm. The first trivial property is that $\tau$ is in $\mathcal{{H}}_{t}(\tau)$ itself:
$\tau\in\mathcal{{H}}_{t}(\tau).$ (21)
The second property is that if $\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)$, then we have
$\mathcal{{H}}_{t}(\tau^{\prime})=\mathcal{{H}}_{t}(\tau).$ (22)
This holds since if $\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)$ and $\tau\in\mathcal{{H}}_{t}(\tau)$ due to (21), $\tau$ and $\tau^{\prime}$ are in the same side from $f_{t}$ (i.e., either $\tau\leq f_{t}\wedge\tau^{\prime}\leq f_{t}$ or $\tau>f_{t}\wedge\tau^{\prime}>f_{t}$), thus $\hat{{S}}(\mathbf{x}_{t};\tau)=\hat{{S}}(\mathbf{x}_{t};\tau^{\prime})$, implying $\mathcal{{H}}_{t}(\tau^{\prime})=\mathcal{{H}}_{t}(\tau)$ by our construction of $\mathcal{{H}}_{t}(\cdot)$ in our algorithm. The second case when $\tau\in\mathcal{{H}}_{t}(\tau^{\prime})$ similarly holds as well.
The third property, similar to the second one, is that if $\tau^{\prime}\in\mathcal{{H}}_{t}^{C}(\tau)$, then we have
$\mathcal{{H}}_{t}^{C}(\tau)=\mathcal{{H}}_{t}(\tau^{\prime}),\text{or equivalently}\mathcal{{H}}_{t}(\tau)=\mathcal{{H}}_{t}^{C}(\tau^{\prime}).$ (23)
This property holds as follows: $\tau^{\prime}\in\mathcal{{H}}_{t}(\tau^{\prime})$ holds due to (21). Suppose $\tau^{\prime}\in\mathcal{{H}}_{t}^{C}(\tau)$, then we have $\tau^{\prime}\notin\mathcal{{H}}_{t}(\tau)$. This along with $\tau^{\prime}\in\mathcal{{H}}_{t}(\tau^{\prime})$ implies $\mathcal{{H}}_{t}(\tau^{\prime})\neq\mathcal{{H}}_{t}(\tau)$. Since $\mathcal{{H}}_{t}(\tau_{t})$ is either $\{\tau\mid\tau\leq f_{t}\}$ or $\{\tau\mid\tau>f_{t}\}$ depending on the value of $\tau_{t}$, $\mathcal{{H}}_{t}(\cdot)$ partitions the space $[0,1]$ into two disjoint sets based on $f_{t}$, where one set is the complement of the other. Therefore, $\mathcal{{H}}_{t}(\tau)\neq\mathcal{{H}}_{t}(\tau^{\prime})$ is equivalent to either $\mathcal{{H}}_{t}^{C}(\tau)=\mathcal{{H}}_{t}(\tau^{\prime})$ or $\mathcal{{H}}_{t}(\tau)=\mathcal{{H}}_{t}^{C}(\tau^{\prime})$.
The final property is that
$\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)\quad\text{if and only if}\quad\tau\in\mathcal{{H}}_{t}(\tau^{\prime}).$ (24)
Suppose that $\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)$ and $\tau^{\prime}\in\mathcal{{H}}_{t}(\tau^{\prime})$, then the above holds as follows:
$\tau\in\mathcal{{H}}_{t}(\tau)=\mathcal{{H}}_{t}(\tau^{\prime}),$
where the inclusion holds due to (21) and the equality holds due to (22) if $\tau^{\prime},\tau\in\mathcal{{H}}_{t}(\tau)$. The reverse holds similarly as well.
Recall that $\ell_{t}(\tau)\in[0,\ell_{\text{max}}]$ and the definition of our loss estimator (8). Then, we observe that this estimator is unbiased as follows:
$\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau,\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))$ $=\sum_{\tau^{\prime}\in\mathcal{{H}}}p_{t}(\tau^{\prime})\left[\frac{\ell_{t}(\tau,\alpha)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau^{\prime}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau^{\prime}}))\right]$
$=\ell_{t}(\tau,\alpha)\sum_{\tau^{\prime}\in\mathcal{{H}}}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau^{\prime}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau^{\prime}}))$
$=\ell_{t}(\tau,\alpha)\sum_{\tau^{\prime}\in\mathcal{{H}}}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau^{\prime})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\mathbbm{1}({\tau^{\prime}\in\mathcal{{H}}_{t}({\tau})})$ (25)
$=\ell_{t}(\tau,\alpha)\sum_{\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau^{\prime})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}$ (26)
$=\ell_{t}(\tau,\alpha)\sum_{\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}$ (27)
$=\ell_{t}(\tau,\alpha)\sum_{\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}\mathbbm{1}(\bar{\tau}\in\mathcal{{H}}_{t}({\tau}))\cdot p_{t}(\bar{\tau})}$ (28)
$=\ell_{t}(\tau,\alpha)\sum_{\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}p_{t}(\bar{\tau})}$
$=\ell_{t}(\tau,\alpha).$ (29)
[…]
Also, to avoid notational clutter, we define
[…]
In the following, we find the lower and upper bound of $\mathbbm{E}\ln\frac{W_{T}}{W_{0}}$ to combine them for the final regret bound, where the expectation is taken over $\tau_{t}\sim p_{t}$ from samples drawn in advance by an oblivious adversary.
The upper bound of potentials.
We find an upper bound of $\mathbbm{E}\ln\frac{W_{T}}{W_{0}}$. To this end, we first consider an upper bound of $\mathbbm{E}\ln\frac{W_{t}}{W_{t-1}}$. In particular, we have
$\ln\frac{W_{t}}{W_{t-1}}$ $=\ln\frac{\sum_{i=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t}^{i})}{\sum_{j=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t-1}^{j})}$
$=\ln\frac{\sum_{i=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t-1}^{i})\,\exp(-\eta\tilde{\ell}_{t}^{i})}{\sum_{j=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t-1}^{j})}$
$=\ln\sum_{i=1}^{|\mathcal{H}|}\frac{\exp(-\eta\tilde{L}_{t-1}^{i})}{\sum_{j=1}^{|\mathcal{H}|}\exp(-\eta\tilde{L}_{t-1}^{j})}\exp(-\eta\tilde{\ell}_{t}^{i})$
$=\ln{\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})}\exp(-\eta\tilde{\ell}_{t}^{i})$ (30)
$\leq\ln\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\left(1-\eta\tilde{\ell}_{t}^{i}+\frac{\eta^{2}(\tilde{\ell}_{t}^{i})^{2}}{2}\right)$ (31)
$=\ln\left(1-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})(\tilde{\ell}_{t}^{i})^{2}\right)$
$\leq-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})(\tilde{\ell}_{t}^{i})^{2}$ (32)
$=-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\left(\frac{\ell_{t}(\tau_{i},\alpha)\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\tau_{t}}))}{{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}}\right)\ell_{t}(\tau_{i},\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))$ (33)
${=}-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{i=1}^{|\mathcal{H}|}\frac{{\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\tau_{t}}))\cdot p_{t}(\tau_{i})}\cdot\ell_{t}(\tau_{i},\alpha)\cdot\ell_{t}(\tau_{i},\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}$
${=}-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{\tau_{i}\in\mathcal{{H}}_{t}({\tau_{t}})}\frac{{p_{t}(\tau_{i})}\cdot\ell_{t}(\tau_{i},\alpha)\cdot\ell_{t}(\tau_{i},\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau_{t})}\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}$
${=}-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{\tau_{i}\in\mathcal{{H}}_{t}(\tau_{t})}\frac{{p_{t}(\tau_{i})}\cdot\ell_{t}(\tau_{i},\alpha)^{2}}{(\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau_{t})}\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau}))^{2}}$
${\leq}-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\ell_{\text{max}}^{2}\frac{\eta^{2}}{2}\sum_{\tau_{i}\in\mathcal{{H}}_{t}({\tau_{t}})}\frac{{p_{t}(\tau_{i})}}{(\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau}))^{2}}$
$=-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\ell_{\text{max}}^{2}\frac{\eta^{2}}{2}\sum_{\tau_{i}\in\mathcal{{H}}_{t}({\tau_{t}})}\frac{{p_{t}(\tau_{i})}}{(\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau_{i}\in\mathcal{{H}}_{t}({\tau_{t}}))\cdot p_{t}(\bar{\tau}))^{2}}$ (34)
$=-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\ell_{\text{max}}^{2}\frac{\eta^{2}}{2}\sum_{\tau_{i}\in\mathcal{{H}}_{t}({\tau_{t}})}\frac{{p_{t}(\tau_{i})}}{(\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau}))^{2}}$ (35)
$=-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\ell_{\text{max}}^{2}\frac{\eta^{2}}{2}\frac{1}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}$
$=-\eta\mathbbm{E}_{\tau_{i}\sim p_{t}}\tilde{\ell}_{t}^{i}+\ell_{\text{max}}^{2}\frac{\eta^{2}}{2}\frac{1}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}.$
Here, (30) follows directly from the definition of the hypotheses distribution $p_{t}(\tau_{t})$, (31) is obtained by applying the second-order Taylor bound $e^{x}\leq 1+x+{x^{2}}/{2}$ which is valid for $x\leq 0$, together with the observation that $-\eta\tilde{\ell}_{t}^{i}\leq 0$, (32) hold by the standard logarithmic bound $\ln(1+x)\leq x$ for any $x\in\mathbb{R}$. (33) follows from the definition of our estimator. (34) holds as $\tau_{i},\bar{\tau}\in\mathcal{{H}}_{t}(\bar{\tau})=\mathcal{{H}}_{t}(\tau_{t})$ due to conditions in summations and (22) i.e., $\tau_{i}\in\mathcal{{H}}_{t}(\tau_{t})$ and $\bar{\tau}\in\mathcal{{H}}_{t}(\tau_{t})$ in the summations derive $\mathcal{{H}}_{t}(\tau_{t})=\mathcal{{H}}_{t}(\tau_{i})=\mathcal{{H}}_{t}(\bar{\tau})$ from (22). Finally, (35) holds as $\tau_{i}\in\mathcal{{H}}_{t}(\tau_{t})$ is always true due to the outer summation condition.
Then, we have the following:
$\mathbbm{E}\left[\ln\frac{W_{T}}{W_{0}}\right]$ $=\mathbbm{E}\left[\sum_{t=1}^{T}\ln\frac{W_{t}}{W_{t-1}}\right]$
$\leq\mathbbm{E}\left[-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{t=1}^{T}\frac{\ell_{\text{max}}^{2}}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}\right]$ (36)
$=-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\frac{\ell_{\text{max}}^{2}}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})},$
$=-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}+{\eta^{2}}\ell_{\text{max}}^{2}T,$ (37)
where, after the telescoping sum, (36) applies the exponential-weights bound obtained in the previous step to each factor $\ln\frac{W_{t}}{W_{t-1}}$ and (37) holds as follows:
$\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}$ $\frac{\ell_{\text{max}}^{2}}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}$
$=\sum_{t=1}^{T}\sum_{\tau_{t}\in\mathcal{{H}}}\frac{p_{t}(\tau_{t})\ell_{\text{max}}^{2}}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}$
$=\sum_{t=1}^{T}\ell_{\text{max}}^{2}\sum_{\tau_{t}\in\mathcal{{H}}}\frac{p_{t}(\tau_{t})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}$
$=\sum_{t=1}^{T}\ell_{\text{max}}^{2}\left(\sum_{\tau\in\mathcal{{H}}_{t}(\tau_{t})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}p_{t}(\bar{\tau})}+\sum_{\tau\in\mathcal{{H}}_{t}^{C}(\tau_{t})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}p_{t}(\bar{\tau})}\right)$
$=\sum_{t=1}^{T}\ell_{\text{max}}^{2}\left(\sum_{\tau\in\mathcal{{H}}_{t}(\tau_{t})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}+\sum_{\tau\in\mathcal{{H}}_{t}^{C}(\tau_{t})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}p_{t}(\bar{\tau})}\right)$ (38)
$=\sum_{t=1}^{T}\ell_{\text{max}}^{2}\left(\sum_{\tau\in\mathcal{{H}}_{t}(\tau_{t})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}+\sum_{\tau\in\mathcal{{H}}_{t}^{C}(\tau_{t})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in{\mathcal{{H}}_{t}^{C}(\tau_{t})}}p_{t}(\bar{\tau})}\right)$ (39)
$={2}\ell_{\text{max}}^{2}T,$ (40)
where (38) holds as (22), i.e., $\mathcal{{H}}_{t}(\tau)=\mathcal{{H}}_{t}(\tau_{t})$ due to $\tau\in\mathcal{{H}}_{t}(\tau_{t})$ in the summation. Also, (39) holds due to (23), i.e., $\mathcal{{H}}_{t}(\tau)=\mathcal{{H}}_{t}^{C}(\tau_{t})$ as $\tau\in\mathcal{{H}}_{t}^{C}(\tau_{t})$ from the summation condition.
The lower bound of potentials.
Then, we find a lower bound of $\mathbbm{E}\left[\ln\frac{W_{T}}{W_{0}}\right]$. In particular, we have
$\mathbbm{E}\left[\ln{\frac{W_{T}}{W_{0}}}\right]$ $=\mathbbm{E}\left[\ln\sum_{i=1}^{|\mathcal{{H}}|}\exp\{-\eta\tilde{L}_{T}^{i}\}\right]-\ln|\mathcal{{H}}|$
$\geq\mathbbm{E}\left[\ln\left(\max_{i\in\{1,\dots,|\mathcal{{H}}|\}}\exp\{-\eta\tilde{L}^{i}_{T}\}\right)\right]-\ln|\mathcal{{H}}|$
$=\mathbbm{E}\left[-\eta\min_{i\in\{1,\dots,|\mathcal{{H}}|\}}\tilde{L}^{i}_{T}\right]-\ln|\mathcal{{H}}|$
$=\mathbbm{E}\left[-\eta\min_{i\in\{1,\dots,|\mathcal{{H}}|\}}\sum_{t=1}^{T}\tilde{\ell}^{i}_{t}\right]-\ln|\mathcal{{H}}|$
$\geq-\eta\min_{i\in\{1,\dots,|\mathcal{{H}}|\}}\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\tilde{\ell}^{i}_{t}-\ln|\mathcal{{H}}|$
$=-\eta\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau,\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))-\ln|\mathcal{{H}}|$
${=}-\eta\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}{\ell_{t}(\tau,\alpha)}-\ln|\mathcal{{H}}|,$
where the last equality holds as the fact of the unbiased loss estimator, i.e., $\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell(\tau_{i},\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))=\ell_{t}(\tau_{i},\alpha)$ from (29).
Combine the lower and upper bounds.
Finally, we combine the lower and upper bound of $\mathbbm{E}\ln\frac{W_{T}}{W_{0}}$ to derive our regret bound. In particular, we have
$\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau,\alpha)$ $\leq{\eta T}\ell_{\text{max}}^{2}+\frac{\ln|\mathcal{{H}}|}{\eta},$
Then, substituting the optimal learning rate $\sqrt{\frac{\ln|\mathcal{H}|}{\ell_{\text{max}}^{2}T}}$, minimizing the upper bound, into the preceding bound yields the claimed regret guarantee, which completes the proof.
Appendix H Discussion on Selection Efficiency
In this section, we demonstrate that our algorithm always converges to a non-trivial selective generator, i.e., $\operatorname*{arg\,min}_{\tau}\sum_{t=1}^{T}[a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)]\notin\mathcal{{H}}^{\text{triv}}$, if there exists some $\tau\in\mathcal{{H}}\setminus\mathcal{{H}}^{\text{triv}}$ that satisfies the FDR guarantee, where $\mathcal{{H}}^{\text{triv}}\coloneqq\{\tau\mid\sum_{t=1}^{T}a_{t}(\tau)=T\}$.
Let $\tau^{\text{triv}}\in\mathcal{{H}}^{\text{triv}}$, which is always abstaining, then for any $\tau^{\text{non-triv}}\in\mathcal{{H}}\setminus\mathcal{{H}}^{\text{triv}}$ that satisfies the FDR guarantee, the following inequality holds,
$\frac{\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq_{E}\mathbf{y}_{t}\right)}{\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\right)}\leq\alpha$
$\Rightarrow\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq_{E}\mathbf{y}_{t}\right)\leq\alpha\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\right).$ (41)
Let $L_{T}(\tau,\alpha)\coloneqq\sum_{t=1}^{T}[a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)]$, which is the cumulative loss of $\tau$, and consider the difference of cumulative loss between $\tau^{\text{triv}}$ and $\tau^{\text{non-triv}}$,
$L_{T}(\tau^{\text{triv}})-$ $L_{T}(\tau^{\text{non-triv}})$
$=\sum_{t=1}^{T}[a_{t}(\tau^{\text{triv}})+\lambda d_{t}(\tau^{\text{triv}},\alpha)]-\sum_{t=1}^{T}[a_{t}(\tau^{\text{non-triv}})+\lambda d_{t}(\tau^{\text{non-triv}},\alpha)]$
$=\sum_{t=1}^{T}\bigg{[}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{triv}})=\texttt{IDK}\right)$
$\qquad\qquad+\lambda\left(\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{triv}})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau^{\text{triv}})\neq_{E}\mathbf{y}_{t})-\alpha\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{triv}})\neq\texttt{IDK})+\alpha\right)\bigg{]}$
$\qquad-\sum_{t=1}^{T}\bigg{[}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})=\texttt{IDK}\right)$
$\qquad\qquad+\lambda\left(\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq_{E}\mathbf{y}_{t})-\alpha\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK})+\alpha\right)\bigg{]}$
$=T+\lambda\alpha T-\Bigg{[}T-\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-non-triv}})\neq\texttt{IDK}\right)$
$\qquad\qquad\qquad+\lambda\bigg{\{}\alpha\bigg{(}T-\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\right)\bigg{)}$
$\qquad\qquad\qquad\qquad+\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq_{E}\mathbf{y}_{t}\right)\bigg{\}}\Bigg{]}$
$=\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\right)+\lambda\alpha\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\right)$
$\qquad\qquad\qquad\qquad-\lambda\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq_{E}\mathbf{y}_{t}\right).$
Since $\tau^{\text{non-triv}}$ satisfies the FDR guarantee (41), we get
$L_{T}(\tau^{\text{triv}})-L_{T}(\tau^{\text{non-triv}})\geq\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\right)=T-\sum_{t=1}^{T}a_{t}(\tau^{\text{non-triv}})>0,$ (42)
where $T-\sum_{t=1}^{T}a_{t}(\tau^{\text{non-triv}})>0$ holds due to $\tau^{\text{non-triv}}\notin\mathcal{{H}}^{\text{triv}}$. This implies that, given sufficiently many data points, any $\tau^{\text{non-triv}}$ that satisfies the FDR guarantee incurs at least $T-\sum_{t=1}^{T}a_{t}(\tau^{\text{non-triv}})$ less loss than $\tau^{\text{triv}}$. This is empirically supported by Figure 21, 22, and 25.